OpenStudy (anonymous):

4. If the measure of the exterior angles of a regular polygon is 60 degrees , how many sides does the polygon have? 5. The angle sum in a regular 5 -gon is _______. A) 180*5 B) 180 * 4 C) 180 * 3 D) None of the other choices

OpenStudy (anonymous):

d

OpenStudy (saifoo.khan):

HOw @ngocthi0101

OpenStudy (anonymous):

yes, explain!?

OpenStudy (anonymous):

a+b+c+d+e = 360

OpenStudy (anonymous):

Recall: for a convex polygon, the sum of the exterior angles will always be 60. If a polygon is regular, all of its exterior angles will be equal. So, m of ext angle in a reg polygon = 360/n where n is the number of sides. Transforming the equation, we can discover that: number of sides/vertices/angles/whatever you wanna call it = 360/m of ea ext angle = 360/60 = 6

OpenStudy (anonymous):

THM: sum of int angles in a convex polygon with n sides is (n - 2)180 Plugging in our value of n which is 5, we get 3*180 = 180*3 so the answer is C.

OpenStudy (anonymous):

would i use this formula? (n-2) × 180°

OpenStudy (anonymous):

thank you, bluepig

OpenStudy (anonymous):

The measure of an exterior angle in a regular 11-gon is _______. The measure of a vertex angle in a regular 15 -gon is ________.

OpenStudy (anonymous):

You know enough now to solve the first question. You know enough to figure out how to solve the second one, but you'll need to think a bit. I think it'll be more rewarding if you do it yourself and you're definitely capable of it.

OpenStudy (anonymous):

checking my answers

OpenStudy (anonymous):

360/11 approx = 32.727 (15 - 2)180/15 = 13*180/15 = 156